OpenStudy (anonymous):

find the zeros of the function algebraically (show work please) f(x)= (square rooted) 2x(end square root) -1

OpenStudy (jamesj):

The function is: \[f(x) = \sqrt{2x} - 1\] ?

OpenStudy (jamesj):

Assuming yes, then \[f(x) = \sqrt{2x} - 1 = 0 \ \ \ \ \ \hbox{ then } \ \ \ \sqrt{2x} = 1\]

OpenStudy (jamesj):

Squaring both sides 2x = 1 So ...